5. (Fully graded)An observer in frame S sees lightning simultaneously strike two points 100 m apart. The first strike occurs at zi = V1 =句= t,-0 and the second at z2 = 100m, y2 = 뀨2 = t2 = 0. (a) What are the coordinates of these two events in a frame S moving at 0.70c in the positive x direction relative to S? b) How far apart are the events in S? (c) Are the events simultaneous in S? If not, what is the difference in time between the events and which event occurs first?
Throughout this homework assignment, you may use e to denote the speed of light in both your work and final answers without using its numerical value. (This is actually preferred unless the problem requires the numerical value of e.)
Answer #1

Frame S' is moving along the common X axis relative to frame S with a speed v=0.70c.

Lorentz transformations relating are

x' =rac{x-vt}{sqrt{1-rac{v^2}{c^2}}}, y'=y, z'=z and C2

(a)

The two events in the S frame are (x_1,y_1,z_1,t_1)=(0,0,0,0) and (22. y2, 32, t2) ( 100 m. 0. 0. 0)

t1 r1- 0.70ct10 - 0.70c 0 VI-0.7021-0.702 C2

y'_1=y_1=0

z'_1=z_1=0

t'_1=rac{t_1-rac{vx_1}{c^2}}{sqrt{1-rac{v^2}{c^2}}}=rac{0-rac{0.70c*0}{c^2}}{sqrt{1-0.70^2}}=0,s

The two events in the S' frame are

(x'_1,y'_1,z'_1,t'_1)=(0,m,0,0,0,s)

x'_2 =rac{x_2-vt_2}{sqrt{1-rac{v^2}{c^2}}}=rac{x_2-0.70ct_2}{sqrt{1-0.70^2}}=rac{100-0.70c*0}{sqrt{1-0.70^2}}=140.028,m

y'_2=y_2=0

z'_2=z_2=0

t'_2=rac{t_2-rac{vx_2}{c^2}}{sqrt{1-rac{v^2}{c^2}}}=rac{0-rac{0.70c*100}{c^2}}{sqrt{1-0.70^2}}=-rac{98.019606}{c}=-3.269582*10^{-7},s

(z, . Љ. 2,, t.) (140.028 m. 0. 0.-3.269582 * 10-7 s)

(b)

The events are 2 1140.028-0 140.028 m apart

(c)

The events are not simultaneous.

t'_1=0,s

t'_2=-3.269582*10^{-7},s

The time difference is t1 -t = 3.269582 * 10-s

The second event (with subscript 2) occurred first.